OpenStudy (anonymous):

What is the sum of three consecutive even integers if m is the middle integer???? Thanks for the help!

OpenStudy (help!!!!):

So if they're consecutive they must add +1

OpenStudy (campbell_st):

the lower is m - 1 and the number above m is m + 1 so the sum is m - 1 + m + m + 1 = hope it helps

OpenStudy (anonymous):

if m is the middle term,then three consecutive even integers are m-2,m,m+2 then sum=m-2+m+m+2=3m